OpenStudy (anonymous):

how do you find the equation for a line that passes through two coordinates (2,1), (5,3)?

OpenStudy (zzr0ck3r):

what is the slope of the line connecting those two points?

OpenStudy (anonymous):

there isn't a slope

OpenStudy (anonymous):

the options are 3y=2x+1 3y=2x-1 3y=-2x-1 3y=-2x+1

OpenStudy (zzr0ck3r):

im not telling you the answer im trying to show you how to get it. what is the slope of the line through the two points listed? y_1-y_0/x_1-x_0

OpenStudy (anonymous):

3/4

OpenStudy (zzr0ck3r):

(2,1) and (5,3) \[\frac{3-1}{5-2}=?\]

OpenStudy (anonymous):

2/3?

terenzreignz (terenzreignz):

Lack of confidence is disturbing :) Given two points, the slope is "difference of the y's divided by difference of the x's" :D

OpenStudy (zzr0ck3r):

yes 2/3 ok so we know it has the form y=mx+b where m is slope so y=2/3x+b right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right

terenzreignz (terenzreignz):

Okay, whenever you have a slope \(\large \color{blue}m\) and a point \(\large (\color{red}p \ , \color{green}q)\) The equation of the line having that slope and passing through said point is given by: \[\Large y = \color{blue}m(x-\color{red}p)+\color{green}q\] You may use any one of your two given points for \(\large (\color{red}p \ , \color{green}q)\)

OpenStudy (anonymous):

none of this makes sense.

terenzreignz (terenzreignz):

Okay... simply put... choose a point from the two that you were given. It doesn't really matter which, just choose one.

OpenStudy (unklerhaukus):

another method you could use is; to plot the points draw the line and determine the equation from the picture 

OpenStudy (unklerhaukus):

another method you could use is; to plot the points and draw the line and determine the equation from the picture 

terenzreignz (terenzreignz):

Well, anyway, once you've chosen from those points, the left coordinate would be \(\large \color{red}p\) and the right coordinate would be \(\large \color{green}q\). You already have your slope \(\large \color{blue}m\). So just plug them into this equation and simplify:\[\Large y = \color{blue}m(x-\color{red}p)+\color{green}q\]